The table below shows the lbs of ice cream sold at a local ice cream parlor for the previous 10 weeks. Assume you are helping the owner develop a weekly forecasting model.

Week lbs of ice cream
1 110
2 122
3 108
4 110
5 121
6 114
7 103
8 112
9 117
10 108

Find the FIT for week 7, using an exponential smoothing with trend model, for the data using alpha of 0.2 along with a delta (δ) of 0.1. Assume an initial trend value (T1) of 1 and an initial forecast (F1) of 105.
Answer #1

Instead of beta, consider delta

For Period 2:

F = 110 + 0.2*(110-105) = 106.00

T = 1 + 0.1*(106 - 105 - 1) = 1.00

TAF = 106 + 1 = 107

Period Actual F T TAF
1 110      105.00          1.00
2 122      106.00          1.00     107.00
3 108      109.20          1.22     110.42
4 110      108.96          1.07     110.03
5 121      109.17          0.99     110.16
6 114      111.53          1.13     112.66
7 103      112.03          1.06     113.09
8 112      110.22          0.78     111.00
9 117      110.58          0.73     111.31
10 108      111.86          0.79     112.65

For period 7, TAF = 113.09